# Sequential estimation of disturbed aerodynamic flows from sparse measurements via a reduced latent space

### v1

@@ 2.2 Finding the low-dimensional latent space
Figure 3: Network architecture in the present study, including a non-linear physics-augmented autoencoder. The observation operator that maps the learned latent vector to the pressure measurements is learned on-the-fly while training the autoencoder.
To enable efficient data assimilation and accurate flow reconstruction, we first construct a compact, low-dimensional representation of the high-dimensional flow state. This reduced-order space is designed to capture the dominant coherent structures in the vorticity field, along with the associated lift dynamics, and serves as the latent state space for sequential filtering. To achieve this, we extend the concept of the lift-augmented autoencoder introduced by Fukami and Taira, (2023) and adapt it to the requirements of our framework. The architecture of the resulting nonlinear physics-augmented autoencoder is illustrated in Figure 3 and table 1. In this model, the high-dimensional vorticity field $\boldsymbol{\omega}\in\mathbb{R}^{l}$ is encoded into a low-dimensional latent vector $\boldsymbol{\xi}\in\mathbb{R}^{n}$ via a neural network encoder $\mathcal{F}_{e}$. The latent representation $\boldsymbol{\xi}$, with $n\ll(l+1)$, is then passed through a decoder $\mathcal{F}_{d}$ to reconstruct not only the original vorticity field but also the corresponding lift coefficient and surface pressure values at prescribed sensor locations. This design ensures that the latent space retains sufficient information relevant to both global flow structure and aerodynamic observables.
@@
Here, the lift and pressure reconstruction terms are incorporated into the loss function $\mathcal{L}_{AE}$ to regularize the learned latent space and ensure that it remains informative with respect to key aerodynamic quantities. Notably, this formulation learns the observation operator—mapping the low-dimensional latent states $\boldsymbol{\xi}$ to surface pressure coefficient measurements, $\boldsymbol{p}=\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\in\mathbb{R}^{d}$—during the training process. Rather than training a separate network to learn the observation mapping post hoc, we integrate it directly into the autoencoder training. This joint learning approach has two key advantages: (1) it constrains the latent space to remain consistent with the pressure signatures, thus promoting sensor-relevant representations; and (2) it improves the accuracy and robustness of the learned observation operator while accelerating convergence during training. By embedding observability constraints (though partial) within the compression task, the resulting latent space not only captures the dominant flow structures but also preserves essential information required for downstream tasks such as state estimation and control. As a result, the decoder reconstructs vorticity field and lift as the original flow states, as well as pressure observations.
@@
Conventional autoencoders process each data snapshot independently, disregarding the temporal correlations inherent in a transient fluid flow. This neglect often results in suboptimal compression, jagged or non-smooth latent trajectories, and poor generalization when extrapolating over time. However, for systems governed by smoothly evolving physical processes, such as transient fluid flow, the reduced latent space should ideally preserve this temporal quality. In other words, smooth evolution in the original state space should be mirrored by equally smooth trajectories in the latent representation to ensure physically meaningful and dynamically consistent encoding. To address this issue, we introduce a temporal regularization term that penalizes abrupt changes in the latent space from one input snapshot to the next in the sequence. Specifically, we enforce smoothness by minimizing the second time derivative of the latent variables. Compared to using a first derivative, this higher-order constraint has been shown to yield more stable and consistent representations during training (Xie et al., 2024). By discouraging sudden accelerations in the latent dynamics, this approach promotes smoother temporal evolution and mitigates spurious jumps, which is particularly important for sequential filtering and forecasting tasks.
@@
The loss weights $\beta_{i}$ in Eq. (2.2), where $i\in\{\boldsymbol{\omega},\boldsymbol{p},C_{L},t\}$, are used to balance the relative contributions of each loss component to the total objective. These weights are carefully tuned to emphasize accurate reconstruction of the vorticity field and surface pressure—both of which are critical for ensuring reliable state estimation via the observation operator—while keeping the lift and temporal smoothness terms at comparable but lower magnitudes. Specifically, the lift and temporal losses are scaled to be approximately one order of magnitude smaller than the dominant reconstruction losses to serve as regularizers without overpowering the main signal. Based on empirical testing, we set $\beta_{\boldsymbol{\omega}}=1.0$, $\beta_{\boldsymbol{p}}=100.0$, $\beta_{C_{L}}=1.0$, and $\beta_{t}=5000.0$. For training, $70\%$ of the total random gust cases—each spanning the full gust–airfoil interaction duration—are used for model optimization, while the remaining $30\%$ are held out for validation and testing. Separating training and test sets by distinct gust encounters—rather than by randomly selecting individual snapshots across all cases—better reflects realistic deployment scenarios. In practice, models are trained on a subset of gust conditions and are expected to generalize to entirely new gust patterns, rather than merely interpolating between seen snapshots from the same cases. Training is terminated when the validation loss converges and plateaus (around epoch $\approx 2000$), indicating stable learning.
@@ 2.3 Sequential filtering
Sequential filtering offers a principled framework for estimating the evolving state probability distribution of a dynamical system by continuously incorporating observational data as it becomes available. This study enables online inference of latent aerodynamic states from sparse and noisy surface pressure measurements, facilitated by learned forecast and observation models discussed in Section 2.4. For a comprehensive overview of sequential filtering methods—including their foundations and applications in aerodynamic contexts—we refer the reader to the recent review by Eldredge and Mousavi, (2025).
@@
Each sequential filtering algorithm step consists of two key substeps: the forecast (or prediction) step and the analysis (or update) step. In the forecast step, the system states—here, the latent variables $\boldsymbol{\xi}$—are advanced in time using a dynamical model $\boldsymbol{f}$. When the governing dynamics are not known a priori, they can be learned from data using a NN, resulting in a learned forecast model $\boldsymbol{f}_{\boldsymbol{W}}$, parameterized by network weights $\boldsymbol{W}$. The formulation assumes that the current state depends only on the immediate past state, thus preserving the Markov property of the underlying system dynamics within the reduced latent space. The forecast step provides an estimate of the prior distribution of states at that instant, $\pi(\boldsymbol{\xi}_{k}|\boldsymbol{p}_{1:k-1})$. Upon the arrival of new measurements, $\boldsymbol{p}_{k}$, the analysis step updates the prior to the posterior distribution, $\pi(\boldsymbol{\xi}_{k}|\boldsymbol{p}_{1:k})$. This Bayesian inference problem requires an observation operator $\boldsymbol{h}$ that maps the latent state $\boldsymbol{\xi}$ to observables (e.g., surface pressures), with additive observation noise $\boldsymbol{\eta}$. As with the forecast model, this observation operator can be learned from data using a NN, $\boldsymbol{h}_{\boldsymbol{W}}$, parameterized by its own set of weights.
@@
In the classical setting, when both process and measurement noises are Gaussian and the forecast and observation operators are linear, the posterior distribution of the system states at each time step remains Gaussian and can be computed analytically. However, in fluid dynamic applications such as the one considered in this study, the forecast and observation operators are highly nonlinear, especially when learned via NNs. As a result, the posterior distribution becomes non-Gaussian, even if the noise distributions remain Gaussian. To address this, various extensions of the standard Kalman filter have been developed for nonlinear systems. Among them, the ensemble Kalman filter (EnKF) has proven particularly effective, especially in high-dimensional state spaces. Originally introduced by Evensen, (1994), the EnKF approximates the distribution of system states using a finite ensemble of samples. It has been successfully applied to a wide range of problems, including in fluid dynamics and reduced-order modeling (da Silva and Colonius, 2018, Moldovan et al., 2021, Le Provost and Eldredge, 2021, Eldredge and Mousavi, 2025).
@@
In the EnKF framework, an ensemble of latent states $\{\boldsymbol{\xi}^{i}\}_{i=1}^{M}$ is drawn from some initial prior distribution. During the forecast step, each ensemble member is independently propagated forward in time using the learned dynamics model:
@@
where $\boldsymbol{w}\sim\mathcal{N}(0,\boldsymbol{Q})$ represents the process noise and $\boldsymbol{Q}$ is an empirical covariance, determined from the residual differences between predicted and true latent trajectories over all test cases. This approach is designed to balance the tradeoff between bias and variance in the ensemble of states.
@@
where $\boldsymbol{\eta}$ denotes the observation noise, modeled as independent and identically (Gaussian) distributed (i.i.d.) random variables with covariance $\boldsymbol{\Sigma}_{\boldsymbol{p}}$. In this context, its standard deviation (square root of diagonal elements in $\boldsymbol{\Sigma}_{\boldsymbol{p}}$) is fixed for all sensors and set equal to the mean observation error computed from the learned observation operator on the test aerodynamic cases (except in some experiments in which the noise of selected sensors is increased to simulate sensor failure). Similar to the process noise, this approach attempts to balance bias and variance in the measurement ensemble.
@@
$\boldsymbol{\xi}^{i}_{k}=\boldsymbol{\xi}^{i}_{k|k-1}+\boldsymbol{K}_{k}\left(\boldsymbol{p}^{*}_{k}-\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi}^{i}_{k|k-1})-\boldsymbol{\eta}^{i}_{k}\right),$ (7)
@@
where $\boldsymbol{p}_{k}^{*}$ denotes the true observation, and the term in parentheses is referred to as the innovation—the discrepancy between the actual and predicted observations. The Kalman gain $\boldsymbol{K}_{k}$ determines how much weight is given to the new observation relative to the predicted state and is computed as $\boldsymbol{K}=\boldsymbol{\Sigma}_{\boldsymbol{\xi}\boldsymbol{p}}\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1}$, where $\boldsymbol{\Sigma}_{\boldsymbol{\xi}\boldsymbol{p}}$ is the cross-covariance between the state and observation ensembles; it is approximated empirically from the ensemble. This update strategy enables EnKF to provide efficient and robust posterior estimation, even in complex, nonlinear, and partially observed systems.
Although the EnKF operates in a low-dimensional latent space, spurious correlations between latent states and observations can still arise, particularly when latent variables encode global flow features or when the system involves elliptic couplings such as the pressure Poisson equation. These correlations may introduce non-physical corrections in directions that are weakly constrained by the sensors. To mitigate this, we adopt the low-rank Ensemble Kalman Filter (LREnKF) of Le Provost et al., (2022), which restricts the update to the dominant eigenmodes of the state–observation cross-covariance. This projection acts as an implicit localization, filtering out poorly observed directions and yielding more efficient, stable updates. The method has been demonstrated successfully in point-vortex estimation problems (Le Provost et al., 2022, Eldredge and Mousavi, 2025) and is well-suited to our setting with limited sensor coverage and complex latent structure.
The LREnKF effectively rotates the state and observation spaces into lower-dimensional subspaces spanned by the most informative directions in each space, and performs the state update in these subspaces. Following Cui and Zahm, (2021), in a nonlinear Gaussian setting, the most informative correction directions in the state space can be identified via the state space Gramian, defined as:
@@
$\boldsymbol{C}_{\boldsymbol{\xi}}=\int\left(\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)^{T}\left(\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)\mathrm{d}\pi(\boldsymbol{\xi})\in\mathbb{R}^{n\times n},$ (8)
@@
where the expectation is taken with respect to the prior distribution. Here, $\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})$ denotes the Jacobian of the observation operator, computed via automatic differentiation within the NN framework, and $\boldsymbol{\Sigma}_{\boldsymbol{\xi}}$ is the prior state covariance matrix. By construction, $\boldsymbol{C}_{\boldsymbol{\xi}}$ is positive semi-definite, and its eigendecomposition $\boldsymbol{C}_{\boldsymbol{\xi}}=\boldsymbol{V}\boldsymbol{\Lambda}_{\boldsymbol{\xi}}^{2}\boldsymbol{V}^{T}$ yields an orthonormal basis $\boldsymbol{V}\in\mathbb{R}^{n\times n}$ for the state space, with corresponding eigenvalues $\boldsymbol{\Lambda}_{\boldsymbol{\xi}}\in\mathbb{R}^{n\times n}$. In a similar manner, the observation space Gramian $\boldsymbol{C}_{\boldsymbol{p}}$ was defined by Le Provost et al., (2022) as
@@
$\boldsymbol{C}_{\boldsymbol{p}}=\int\left(\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)\left(\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)^{T}\mathrm{d}\pi(\boldsymbol{\xi})\in\mathbb{R}^{d\times d}.$ (9)
Like its state space counterpart, $\boldsymbol{C}_{\boldsymbol{p}}$ is positive semi-definite, and its eigendecomposition $\boldsymbol{C}_{\boldsymbol{p}}=\boldsymbol{U}\boldsymbol{\Lambda}_{\boldsymbol{p}}^{2}\boldsymbol{U}^{T}$ yields an orthonormal basis $\boldsymbol{U}\in\mathbb{R}^{d\times d}$ for the observation space, with corresponding eigenvalues $\boldsymbol{\Lambda}_{\boldsymbol{p}}\in\mathbb{R}^{d\times d}$. Both Gramians can be efficiently approximated via the Monte Carlo method over the prior ensemble. In practice, the eigenvalues are sorted in descending order, and only the leading $r_{\boldsymbol{\xi}}\leq n$ eigenmodes of the state space Gramian and the leading $r_{\boldsymbol{p}}\leq d$ eigenmodes of the observation space Gramian are retained for assimilation. These ranks are typically chosen to capture a prescribed fraction $\in[0,1]$ of the cumulative normalized energy in the corresponding eigenvalue spectra.
The Kalman update can then be performed efficiently within this informative subspace by projecting the prior states onto the dominant eigenvectors $\boldsymbol{V}$ in the state space and projecting the innovation onto the dominant eigenvectors $\boldsymbol{U}$ in the observation space. After correction in this reduced subspace, the updated states are lifted back into the original latent space at each assimilation step. Accordingly, the Kalman update formula in Eq. (7) can be factorized in the LREnKF framework as:
$\boldsymbol{\xi}^{i}_{k}=\boldsymbol{\xi}^{i}_{k|k-1}+\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\boldsymbol{V}\breve{\boldsymbol{K}}_{k}\boldsymbol{U}^{T}\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1/2}\left(\boldsymbol{p}^{*}_{k}-\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi}^{i}_{k|k-1})-\boldsymbol{\eta}^{i}_{k}\right),$ (10)
where the reduced Kalman gain is computed in the low-dimensional informative subspace as $\breve{\boldsymbol{K}}=\boldsymbol{\Sigma}_{\breve{\boldsymbol{\xi}}\breve{\boldsymbol{p}}}\boldsymbol{\Sigma}_{\breve{\boldsymbol{p}}}^{-1}\in\mathbb{R}^{r_{\boldsymbol{\xi}}\times r_{\boldsymbol{p}}}$. Here, $\boldsymbol{\Sigma}_{\breve{\boldsymbol{\xi}}\breve{\boldsymbol{p}}}$ and $\boldsymbol{\Sigma}_{\breve{\boldsymbol{p}}}$ denote the state-observation cross-covariance and observation covariance, respectively, computed from the projected ensemble members. During each state correction, the innovation for ensemble member $i$ is first computed from the terms inside the parentheses, representing the difference between the measured and predicted pressures at the sensor locations. This innovation is then whitened by multiplying by the inverse square root of the observation error covariance, $\boldsymbol{\Sigma}_{\boldsymbol{p}}^{-1/2}$, to normalize the observation noise to unit variance. The whitened innovation is subsequently projected onto a reduced-dimensional observation subspace spanned by the dominant observation modes, $\boldsymbol{U}$. Within this subspace, the reduced Kalman gain, $\breve{\boldsymbol{K}}_{k}$, maps the informative observation directions to the corresponding directions in the state subspace. The resulting update is then lifted back to the full state space using the dominant state modes, $\boldsymbol{V}$, and scaled (de-whitened) by the square root of the state covariance, $\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}$. A detailed description of the LREnKF algorithm is provided in Le Provost et al., (2021). This formulation ensures that the assimilation update only acts along directions that are both dynamically significant and well-informed by the available measurements, thereby improving robustness and reducing the impact of spurious correlations.

### v2

@@ 2.2 Finding the low-dimensional latent space
Figure 3: Network architecture in the present study, including a non-linear physics-augmented convolutional autoencoder. The observation operator that maps the learned latent vector to the pressure measurements is learned on-the-fly while training the autoencoder.
To enable efficient data assimilation and accurate flow reconstruction, we first construct a compact, low-dimensional representation of the high-dimensional flow state. This reduced-order space is designed to capture the dominant coherent structures in the vorticity field, along with the associated lift dynamics, and serves as the latent state space for sequential filtering. To achieve this, we extend the concept of the lift-augmented autoencoder introduced by Fukami and Taira [15] and adapt it to the requirements of our framework. The architecture of the resulting nonlinear physics-augmented convolutional autoencoder is illustrated in Figure 3 and table 1. In this model, the high-dimensional vorticity field $\boldsymbol{\omega}\in\mathbb{R}^{l}$ is encoded into a low-dimensional latent vector $\boldsymbol{\xi}\in\mathbb{R}^{n}$ via a NN encoder $\mathcal{F}_{e}$. The latent representation $\boldsymbol{\xi}$, with $n\ll l$, is then passed through a decoder $\mathcal{F}_{d}$ to reconstruct not only the original vorticity field but also the corresponding lift coefficient and surface pressure values at prescribed sensor locations. This design ensures that the latent space retains sufficient information relevant to both global flow structure and aerodynamic observables.
@@
Here, the lift and pressure reconstruction terms are incorporated into the loss function $\mathcal{L}_{AE}$ to regularize the learned latent space and ensure that it remains informative with respect to key aerodynamic quantities. Notably, this formulation learns the observation operator—mapping the low-dimensional latent states $\boldsymbol{\xi}$ to surface pressure coefficient measurements, $\boldsymbol{p}=\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\in\mathbb{R}^{d}$, with nonlinear activation function (tangent hyperbolic)—during the training process. Rather than training a separate network to learn the observation mapping post hoc, we integrate it directly into the autoencoder training. This joint learning approach has two key advantages: (1) it constrains the latent space to remain consistent with the pressure signatures, thus promoting sensor-relevant representations; and (2) it improves the accuracy and robustness of the learned observation operator while accelerating convergence during training. By embedding observability constraints (though partial) within the compression task, the resulting latent space not only captures the dominant flow structures but also preserves essential information required for downstream tasks such as state estimation and control. As a result, the decoder reconstructs the vorticity field and lift as the original flow states, as well as pressure observations.
@@
Conventional autoencoders process each data snapshot independently, disregarding the temporal correlations inherent in a transient fluid flow. As a result, the learned latent manifold may not form a dynamically smooth or physically consistent phase space, leading to suboptimal compression and reduced suitability for learning time-evolving dynamics. However, for systems governed by smoothly evolving physical processes, such as transient fluid flow, the reduced latent space should ideally preserve this temporal quality. In other words, smooth evolution in the original state space should be mirrored by equally smooth trajectories in the latent representation to ensure physically meaningful and dynamically consistent encoding. To address this issue, we introduce a temporal regularization term that penalizes abrupt changes in the latent space from one input snapshot to the next in the sequence. Specifically, we enforce smoothness by minimizing the second time derivative of the latent variables. Compared to using a first derivative, this higher-order constraint has been shown to yield more stable and consistent representations during training [44]. By discouraging sudden accelerations in the latent dynamics, this approach promotes smoother temporal evolution and mitigates spurious jumps, which is particularly important for sequential filtering and forecasting tasks.
@@
The loss weights $\beta_{i}$ in Eq. (2.2), where $i\in\{\boldsymbol{\omega},\boldsymbol{p},C_{L},t\}$, are used to balance the relative contributions of each loss component to the total objective. These weights are carefully tuned to emphasize accurate reconstruction of the vorticity field and surface pressure—both of which are critical for ensuring reliable state estimation via the observation operator—while keeping the lift and temporal smoothness terms at comparable but lower magnitudes. Specifically, the lift and temporal losses are scaled to be approximately one order of magnitude smaller than the dominant reconstruction losses to serve as regularizers without overpowering the main signal. Based on empirical testing, we set $\beta_{\boldsymbol{\omega}}=1.0$, $\beta_{\boldsymbol{p}}=100.0$, $\beta_{C_{L}}=1.0$, and $\beta_{t}=5000.0$. It is important to note that the relatively large coefficient $\beta_{t}$ does not indicate that the temporal regularization term dominates the overall objective. In practice, the discrete second-order temporal increments, $\boldsymbol{\xi}_{t+1}-2\boldsymbol{\xi}_{t}+\boldsymbol{\xi}_{t-1}||_{2}^{2}$, are several orders of magnitude smaller than the reconstruction losses. As a result, a larger weighting factor is required to balance their contribution during training. This choice of scaling allows the temporal smoothness constraint to meaningfully influence the optimization process while maintaining the primacy of the reconstruction objectives.
For training, $70\%$ of the total random gust cases—each spanning the full gust–airfoil interaction duration—are used for model optimization, while the remaining $30\%$ are held out for validation and testing. Separating training and test sets by distinct gust encounters—rather than by randomly selecting individual snapshots across all cases—better reflects realistic deployment scenarios. In practice, models are trained on a subset of gust conditions and are expected to generalize to entirely new gust patterns. Training is terminated when the validation loss converges and plateaus (around epoch $\approx 2000$), indicating stable learning.
@@ 2.3 Sequential filtering
Sequential filtering offers a principled framework for estimating the evolving state probability distribution of a dynamical system by continuously incorporating streaming observational data. This study enables online inference of latent aerodynamic states from sparse and noisy surface pressure measurements, facilitated by learned forecast and observation models discussed in Section 2.4. For a comprehensive overview of sequential filtering methods—including their foundations and applications in aerodynamic contexts—we refer the reader to the recent work by Eldredge and Mousavi [10].
@@
Each sequential filtering algorithm consists of two key substeps: the forecast (or prediction) step and the analysis (or update) step. In the forecast step, the system states—here, the latent variables $\boldsymbol{\xi}$—are advanced in time using a dynamical model $\boldsymbol{f}$. When the governing dynamics are not known a priori, they can be learned from data using a NN, resulting in a learned forecast model $\boldsymbol{f}_{\boldsymbol{W}}$, parameterized by network weights $\boldsymbol{W}$. The formulation assumes that the current state depends only on the immediate past state, thus preserving the Markov property of the underlying system dynamics within the reduced latent space. In the exact Bayesian framework, the forecast step provides an estimate of the prior distribution of states at that instant, $\pi(\boldsymbol{\xi}_{k}|\boldsymbol{p}_{1:k-1})$. Upon the arrival of new measurements, $\boldsymbol{p}_{k}$, the analysis step updates the prior to the posterior distribution, $\pi(\boldsymbol{\xi}_{k}|\boldsymbol{p}_{1:k})$. This Bayesian inference problem requires an observation operator $\boldsymbol{h}$ that maps the latent state $\boldsymbol{\xi}$ to observables (e.g., surface pressures), with additive observation noise $\boldsymbol{\eta}$. As with the forecast model, this observation operator can be learned from data using a NN, $\boldsymbol{h}_{\boldsymbol{W}}$, parameterized by its own set of weights.
@@
In the classical setting, when both process and measurement noises are Gaussian and the forecast and observation operators are linear, the exact posterior distribution of the system states at each time step remains Gaussian and can be computed analytically. However, in aerodynamic applications such as the one considered in this study, the forecast and observation operators are highly nonlinear, including when they are learned through NNs with nonlinear activations. As a result, the posterior distribution becomes non-Gaussian, even if the noise distributions remain Gaussian. To address this, various extensions of the standard Kalman filter have been developed for nonlinear systems. Among them, the EnKF has proven particularly effective, especially in moderate- to high-dimensional spaces. Originally introduced by Evensen [12], the EnKF approximates the distribution of system states, $\pi(\boldsymbol{\xi}_{k}|\boldsymbol{p}_{1:k})$, using a finite ensemble of samples. It has been successfully applied to a wide range of problems, including in fluid dynamics and reduced-order modeling [10, 8, 30, 27].
@@
In the EnKF framework, an ensemble of latent states $\{\boldsymbol{\xi}^{i}\}_{i=1}^{M}$ is drawn from some initial distribution. During the forecast step, each ensemble member is independently propagated forward in time using the learned dynamical model:
@@
to give a prior ensemble of states $\boldsymbol{\Xi}_{k|k-1}\coloneq\{\boldsymbol{\xi}^{i}_{k|k-1}\}_{i=1}^{M}$. In this equation $\boldsymbol{w}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{Q})$ represents the process noise and $\boldsymbol{Q}$ is an empirical covariance, determined from the one-step residual differences between predicted and true latent trajectories averaged over time for all test cases. This approach is designed to balance the tradeoff between bias and variance in the ensemble of states.
@@
where $\boldsymbol{\eta}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\boldsymbol{\eta}})$ denotes the observation noise with covariance $\boldsymbol{\Sigma}_{\boldsymbol{\eta}}$. In this context, its standard deviation (square root of diagonal elements in $\boldsymbol{\Sigma}_{\boldsymbol{\eta}}$) is fixed for all sensors and set equal to the mean observation error computed from the learned observation operator on the test cases (except in some experiments in which the noise of selected sensors is increased to simulate sensor failure). This choice models the observation uncertainty as the combined effect of sensor noise and observation surrogate approximation error. Collectively, the ensemble of noisy predicted observations is denoted by $\boldsymbol{P}_{k}\coloneq\{\boldsymbol{p}^{i}_{k}\}_{i=1}^{M}$.
@@
$\boldsymbol{\xi}^{i}_{k}=\boldsymbol{\xi}^{i}_{k|k-1}+\boldsymbol{K}_{k}\left(\boldsymbol{p}^{*}_{k}-\boldsymbol{p}_{k}^{i}\right),$ (6)
@@
where $\boldsymbol{p}_{k}^{*}$ denotes the synthetically generated noisy measurements obtained from CFD data by adding an i.i.d. random noise term $\boldsymbol{\eta}_{k}$. The term in parentheses is referred to as the innovation—the discrepancy between the measured and predicted observations. The Kalman gain $\boldsymbol{K}_{k}$ determines how much weight is given to the new observation relative to the predicted state and is computed as $\boldsymbol{K}=\boldsymbol{\Sigma}_{\boldsymbol{\xi}\boldsymbol{p}}\,\boldsymbol{\Sigma}_{\boldsymbol{p}\boldsymbol{p}}^{-1}$, where $\boldsymbol{\Sigma}_{\boldsymbol{\xi}\boldsymbol{p}}$ denotes the cross-covariance between the state and observation ensembles, and $\boldsymbol{\Sigma}_{\boldsymbol{p}\boldsymbol{p}}$ represents the covariance of the noise-injected predicted observations [1]. These covariances are approximated empirically from the ensemble in accordance with Asch et al. [1], with details provided in Appendix A. This update strategy enables the EnKF to provide an efficient and robust posterior estimate, even in complex, nonlinear, and partially observed systems.
At each assimilation step, the dominant directions along which measurements correct the latent state are identified using Gramian operators constructed from the covariance-weighted Jacobian of the observation operator, $\nabla\boldsymbol{h}_{\boldsymbol{W}}$. The eigenvectors of the resulting state-space and observation-space Gramians characterize the dominant correction directions during the analysis step of the EnKF [26]. Following Cui and Zahm [7], in a nonlinear Gaussian setting, the correction directions in the state space can be identified via the state-space Gramian, defined as:
@@
$\boldsymbol{C}_{\boldsymbol{\xi}}=\int\left(\boldsymbol{\Sigma}_{\boldsymbol{\eta}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)^{\top}\left(\boldsymbol{\Sigma}_{\boldsymbol{\eta}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)\mathrm{d}\pi(\boldsymbol{\xi})\in\mathbb{R}^{n\times n},$ (7)
@@
where the expectation is taken with respect to the prior distribution. Here, $\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})$ denotes the Jacobian of the observation operator, computed via automatic differentiation within the NN framework, and $\boldsymbol{\Sigma}_{\boldsymbol{\xi}}$ is the prior state covariance matrix. By construction, $\boldsymbol{C}_{\boldsymbol{\xi}}$ is positive semi-definite, and its eigendecomposition $\boldsymbol{C}_{\boldsymbol{\xi}}=\boldsymbol{V}\boldsymbol{\Lambda}_{\boldsymbol{\xi}}^{2}\boldsymbol{V}^{\top}$ yields an orthonormal basis $\boldsymbol{V}\in\mathbb{R}^{n\times n}$ for the state space, with corresponding eigenvalues $\boldsymbol{\Lambda}_{\boldsymbol{\xi}}\in\mathbb{R}^{n\times n}$. In a similar manner, the observation-space Gramian $\boldsymbol{C}_{\boldsymbol{p}}$ was defined by Le Provost et al. [25] as
@@
$\boldsymbol{C}_{\boldsymbol{p}}=\int\left(\boldsymbol{\Sigma}_{\boldsymbol{\eta}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)\left(\boldsymbol{\Sigma}_{\boldsymbol{\eta}}^{-1/2}\nabla\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})\boldsymbol{\Sigma}_{\boldsymbol{\xi}}^{1/2}\right)^{\top}\mathrm{d}\pi(\boldsymbol{\xi})\in\mathbb{R}^{d\times d}.$ (8)
Like its state space counterpart, $\boldsymbol{C}_{\boldsymbol{p}}$ is positive semi-definite, and its eigendecomposition $\boldsymbol{C}_{\boldsymbol{p}}=\boldsymbol{U}\boldsymbol{\Lambda}_{\boldsymbol{p}}^{2}\boldsymbol{U}^{\top}$ yields an orthonormal basis $\boldsymbol{U}\in\mathbb{R}^{d\times d}$ for the observation space, with corresponding eigenvalues $\boldsymbol{\Lambda}_{\boldsymbol{p}}\in\mathbb{R}^{d\times d}$. Both Gramians can be efficiently approximated via the Monte Carlo method over the prior ensemble. In practice, these eigenvalues are sorted in descending order, so the leading eigenvectors in $\boldsymbol{U}$ correspond to the most informative measurement directions. These directions contribute most strongly to correcting the corresponding leading latent eigendirections in $\boldsymbol{V}$. These dominant directions enable us to analyze the informativeness of the sensors during the filtering update step. It should be noted that the trailing (small eigenvalue) state and observation modes, although contributing minimally to the measurement update step, may still play an important role in reconstructing the high-dimensional vorticity field, since all latent variables participate in the decoding process.